# LiveCLKTBench: Towards Reliable Evaluation of Cross-Lingual Knowledge Transfer in Multilingual LLMs

### v2

@@ Abstract
Evaluating cross-lingual knowledge transfer in large language models (LLMs) is challenging, as correct answers in a target language may arise either from genuine transfer or from prior exposure during pre-training. We present LiveCLKTBench, an automated generation pipeline specifically designed to isolate and measure cross-lingual knowledge transfer. Our pipeline identifies self-contained, time-sensitive knowledge entities from real-world domains, filters them based on temporal occurrence, and verifies them against the model’s knowledge. The documents of these valid entities are then used to generate factual questions, which are translated into multiple languages to evaluate transferability across linguistic boundaries. Using LiveCLKTBench, we evaluate several LLMs across five languages and observe that cross-lingual transfer is strongly influenced by linguistic distance and often asymmetric across language directions. While larger models improve transfer, the gains diminish with scale and vary across domains. These findings provide new insights into multilingual transfer and demonstrate the value of LiveCLKTBench as a reliable benchmark for future research.
@@ 1 Introduction
As large language models (LLMs) continue to grow in scale and capability, a central question arises: How can they serve users globally and equitably? Ideally, an LLM should be able to transfer knowledge acquired in one language to others, rather than relearning the same facts separately across languages — a property known as cross-lingual knowledge transfer (pmlr-v119-hu20b; lauscher-etal-2020-zero; hedderich-etal-2020-transfer).
@@
(1) Leakage-Free Evaluation. A major obstacle in evaluating cross-lingual knowledge transfer is contamination and data leakage (ahuja2024contaminationreportmultilingualbenchmarks). When pretrained corpora already include the same knowledge in multiple languages, models can appear to “transfer” information they have merely memorized. A leakage-free benchmark must therefore ensure that the tested knowledge is unseen in the target language, so that correct answers truly reflect transfer.
(2) Grounding in Real-World Knowledge. Some benchmarks employ synthetic or fictitious data to control knowledge injection and reduce contamination (maini2024tofutaskfictitiousunlearning; maheshwari2024efficacysyntheticdatabenchmark). While this strategy offers strong controllability, recent works have shown that fabricated facts can conflict with pre-existing knowledge and degrade performance (wu2024akewassessingknowledgeediting; jan2025datadopingtrueintelligence; chen-etal-2024-unveiling-flaws). Consequently, when a model fails to answer a fictitious question in a particular language, the failure may not stem from a lack of transferability, but rather from the artificial knowledge being too weakly or even adversarially connected to existing knowledge to support meaningful transfer. We therefore argue that benchmarks grounded in real documents better reflect practical scenarios, where models must absorb genuine knowledge and transfer it across languages while leveraging their existing background knowledge.
(3) Frequent Knowledge Update. As LLMs evolve rapidly, static benchmarks become outdated: newer models may already encode most benchmark facts, causing score saturation and reducing diagnostic value  [^1] https://r0bk.github.io/killedbyllm/. Recent initiatives such as RealTimeQA (kasai2024realtimeqawhatsanswer) and LiveBench (white2025livebenchchallengingcontaminationlimitedllm) highlight the importance of designing benchmarks that are regularly refreshed with new data to prevent staleness.
@@ 2.1 Multilingual Benchmarks
A variety of multilingual benchmarks have been proposed to evaluate model performance across many languages (wu2025bitterlessonlearned2000). Some focus on general natural-language tasks such as classification, question answering, or summarization (e.g., XTREME (hu2020xtreme), XGLUE (liang2020xglue), MEGA (ahuja2023megamultilingualevaluationgenerative)), while others emphasize factual knowledge access in multiple languages, such as M3Exam (zhang2023m3exammultilingualmultimodalmultilevel), AGI-Eval (zhong2023agievalhumancentricbenchmarkevaluating), and Global-MMLU (singh2025globalmmluunderstandingaddressing). These benchmarks provide broad coverage of multilingual capabilities but do not directly isolate cross-lingual transfer.
@@ 2.2 Cross-lingual Transfer Benchmarks
Cross-lingual transfer benchmarks test whether knowledge or skills learned in one language generalize to others. Here, skill transfer refers to the generalization of abilities such as summarization or instruction following to unseen languages (shaham-etal-2024-multilingual; chai2024xcotcrosslingualinstructiontuning; asai2023buffetbenchmarkinglargelanguage), whereas knowledge transfer evaluates the reproduction of factual information acquired in one language when queried in another.
Recent studies analyze the unique challenges of cross-lingual knowledge transfer, such as language transfer asymmetry and knowledge representation barriers (rajaee2024analyzingevaluationcrosslingualknowledge; chua2025crosslingualcapabilitiesknowledgebarriers; litschko2025crossdialectinformationretrievalinformation; yao2024datacontaminationcrosslanguage). Complementing these analyses, several dedicated benchmarks have been proposed. For example, ECLeKTic (goldman2025eclektic) benchmarks cross-lingual knowledge transfer by using Wikipedia articles common in one language but rare in others, ensuring answers reflect genuine transfer. Other benchmarks focus on multilingual knowledge editing, evaluating whether updates introduced in one language propagate to others (nie2025bmike53investigatingcrosslingualknowledge; wei2025mlakemultilingualknowledgeediting).
Another important dimension is the evaluation protocol. Some benchmarks use zero-shot evaluation, querying models directly in target languages (malkin-etal-2022-balanced; goldman2025eclektic), while others adopt a fine-tune–then-test approach, where models are trained in one language and tested in others (shaham-etal-2024-multilingual). Work also distinguishes between parametric knowledge, stored in model parameters (goldman2025eclektic; rajaee2024analyzingevaluationcrosslingualknowledge), and contextual knowledge, provided through in-context examples at inference time (mondshine2025englishimpactprompttranslation; asai2021xorqacrosslingualopenretrieval).
@@ 2.3 Relation to Prior Work
Among prior efforts, ECLeKTic (goldman2025eclektic) is most closely related to our work, as it evaluates factual transfer across languages while reducing leakage by selecting documents that are more common in one language than another. While this strategy offers partial control, leakage cannot be fully ruled out, and reliance on Wikipedia limits coverage of dynamic knowledge. In contrast, LiveCLKTBench injects new knowledge via post-training on a single source language, so correct answers in other languages must arise from genuine transfer. By sourcing documents from rapidly updating domains and allowing configurable time windows, it produces more diverse and extensible benchmarks that stay relevant as models evolve.
@@ 3.2 Evaluation Protocol
    Strategy Comparison: For a single model, test different post-training methods (e.g., continual pretraining (ke2023continualpretraininglanguagemodels), supervised fine-tuning (mecklenburg2024injectingnewknowledgelarge), knowledge editing (wu2024akewassessingknowledgeediting)) to identify approaches that yield better transferability.
@@ 3.3 Evaluation Metrics
To quantify cross-lingual transferability, we follow the metric definitions introduced by goldman2025eclektic. Let $\mathcal{L}$ denote the set of evaluation languages. For each ordered pair $(L_{\text{train}},L_{\text{test}})$ with $L_{\text{train}}\neq L_{\text{test}}$, model predictions can be categorized using the contingency matrix in Table 1.
@@ 4 Demonstrating LiveCLKTBench: A Case Study
We evaluate several open-source LLMs using a single knowledge injection strategy: Continual Pre-Training (ke2023continualpretraininglanguagemodels), where models continue pretraining directly on the target documents using the standard causal language modeling objective (next-token prediction) applied to unmodified text. Although this case study highlights Model Comparison, the framework is agnostic to the post-training method and also supports Strategy Comparison, enabling evaluation of different knowledge injection techniques for a single model.
@@
We consider instruction-tuned, open-source models in the 7–9B parameter range, including Gemma  [^3] https://huggingface.co/google/gemma-2-9b-it, Mistral  [^4] https://huggingface.co/mistralai/Ministral-8B-Instruct-2410555https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, Qwen  [^6] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Llama  [^7] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, and OLMo  [^8] https://huggingface.co/allenai/OLMo-2-1124-7B-Instruct. These model families differ in pretraining data, tokenizer design, and alignment strategies, providing a diverse perspective on cross-lingual transfer.
@@ Appendix B Data Quality Verification
To assess the reliability of our benchmark data, we recruited three annotators (two undergraduate and one graduate student) to manually verify the quality of sampled question–answer pairs. Specifically, we randomly sampled 20% of the generated document–question pairs from each domain of the LiveCLKTBench configuration on time-range: January1,2025 and August31,2025.
Each annotator evaluated whether a question–answer pair violated any of the following criteria:
  • The question is unrelated to the source document.
  • The question is relevant but overly general and not specific to the entity.
  • The correctness of the answer cannot be verified from the source document.
  • The answer is verifiable from the source document, but the provided answer is incorrect.
We aggregated the annotations using majority vote and calculated the proportion of correctly verified pairs as the final precision metric. Across all domains, the manual verification yielded an average precision of 95%, indicating that our pipeline consistently generates data with high factual accuracy and strong contextual alignment with the source documents.
@@ Appendix D Entity Document Examples
Soccer Document Template Sports: {sports} League: {league} Match: {home_team} vs {away_team} Date: {date} Score: {home_score} - {away_score} Match Stats ({home_team} vs {away_team}): {stats_block}

### v3

@@ Abstract
Evaluating cross-lingual knowledge transfer in large language models (LLMs) is challenging, as correct answers in a target language may arise either from genuine transfer or from prior exposure during pre-training. We present LiveCLKTBench, an automated generation pipeline specifically designed to isolate and measure cross-lingual knowledge transfer. Our pipeline identifies self-contained, time-sensitive knowledge entities from real-world domains, filters them based on temporal occurrence, and verifies them against the model’s knowledge. The documents of these valid entities are then used to generate factual questions, which are translated into multiple languages to evaluate transferability across linguistic boundaries. Using LiveCLKTBench, we evaluate several LLMs across five languages and observe that cross-lingual transfer is strongly influenced by linguistic distance and often asymmetric across language directions. While larger models improve transfer, the gains diminish with scale and vary across domains. These findings provide new insights into multilingual transfer and demonstrate the value of LiveCLKTBench as a reliable benchmark for future research. [^1] Code and data are available at link.
@@ 1 Introduction
As large language models (LLMs) continue to grow in scale and capability, a central question arises: How can they serve users globally and equitably? Ideally, an LLM should be able to transfer knowledge acquired in one language to others, rather than relearning the same facts separately across languages — a property known as cross-lingual knowledge transfer (Hu et al., 2020b; Lauscher et al., 2020; Hedderich et al., 2020).
@@
(1) Leakage-Free Evaluation. A major obstacle in evaluating cross-lingual knowledge transfer is contamination and data leakage (Ahuja et al., 2024). When pretrained corpora already include the same knowledge in multiple languages, models can appear to “transfer” information they have merely memorized. A leakage-free benchmark must therefore ensure that the tested knowledge is unseen in the target language, so that correct answers truly reflect transfer.
(2) Grounding in Real-World Knowledge. Some benchmarks employ synthetic or fictitious data to control knowledge injection and reduce contamination (Maini et al., 2024; Maheshwari et al., 2024). While this strategy offers strong controllability, recent works have shown that fabricated facts can conflict with pre-existing knowledge and degrade performance (Wu et al., 2024; Jan et al., 2025; Chen et al., 2024). Consequently, when a model fails to answer a fictitious question in a particular language, the failure may not stem from a lack of transferability, but rather from the artificial knowledge being too weakly or even adversarially connected to existing knowledge to support meaningful transfer. We therefore argue that benchmarks grounded in real documents better reflect practical scenarios, where models must absorb genuine knowledge and transfer it across languages while leveraging their existing background knowledge.
(3) Frequent Knowledge Update. As LLMs evolve rapidly, static benchmarks become outdated: newer models may already encode most benchmark facts, causing score saturation and reducing diagnostic value  [^2] https://r0bk.github.io/killedbyllm/. Recent initiatives such as RealTimeQA (Kasai et al., 2024) and LiveBench (White et al., 2025) highlight the importance of designing benchmarks that are regularly refreshed with new data to prevent staleness.
@@ 2.1 Multilingual Benchmarks
A variety of multilingual benchmarks have been proposed to evaluate model performance across many languages (Wu et al., 2025). Some focus on general natural-language tasks such as classification, question answering, or summarization (e.g., XTREME (Hu et al., 2020a), XGLUE (Liang et al., 2020), MEGA (Ahuja et al., 2023)), while others emphasize factual knowledge access in multiple languages, such as M3Exam (Zhang et al., 2023), AGI-Eval (Zhong et al., 2023), and Global-MMLU (Singh et al., 2025). These benchmarks provide broad coverage of multilingual capabilities but do not directly isolate cross-lingual transfer.
@@ 2.2 Cross-lingual Transfer Benchmarks
Cross-lingual transfer benchmarks test whether knowledge or skills learned in one language generalize to others. Here, skill transfer refers to the generalization of abilities such as summarization or instruction following to unseen languages (Shaham et al., 2024; Chai et al., 2024; Asai et al., 2023), whereas knowledge transfer evaluates the reproduction of factual information acquired in one language when queried in another.
Recent studies analyze the unique challenges of cross-lingual knowledge transfer, such as language transfer asymmetry and knowledge representation barriers (Rajaee and Monz, 2024; Chua et al., 2025; Litschko et al., 2025; Yao et al., 2024). Complementing these analyses, several dedicated benchmarks have been proposed. For example, ECLeKTic (Goldman et al., 2025) benchmarks cross-lingual knowledge transfer by using Wikipedia articles common in one language but rare in others, ensuring answers reflect genuine transfer. Other benchmarks focus on multilingual knowledge editing, evaluating whether updates introduced in one language propagate to others (Nie et al., 2025; Wei et al., 2025).
Another important dimension is the evaluation protocol. Some benchmarks use zero-shot evaluation, querying models directly in target languages (Malkin et al., 2022; Goldman et al., 2025), while others adopt a fine-tune–then-test approach, where models are trained in one language and tested in others (Shaham et al., 2024). Work also distinguishes between parametric knowledge, stored in model parameters (Goldman et al., 2025; Rajaee and Monz, 2024), and contextual knowledge, provided through in-context examples at inference time (Mondshine et al., 2025; Asai et al., 2021).
@@ 2.3 Relation to Prior Work
Among prior efforts, ECLeKTic (Goldman et al., 2025) is most closely related to our work, as it evaluates factual transfer across languages while reducing leakage by selecting documents that are more common in one language than another. While this strategy offers partial control, leakage cannot be fully ruled out, and reliance on Wikipedia limits coverage of dynamic knowledge. In contrast, LiveCLKTBench injects new knowledge via post-training on a single source language, so correct answers in other languages must arise from genuine transfer. By sourcing documents from rapidly updating domains and allowing configurable time windows, it produces more diverse and extensible benchmarks that stay relevant as models evolve.
@@ 3.2 Evaluation Protocol
    Strategy Comparison: For a single model, test different post-training methods (e.g., continual pretraining (Ke et al., 2023), supervised fine-tuning (Mecklenburg et al., 2024), knowledge editing (Wu et al., 2024)) to identify approaches that yield better transferability.
@@ 3.3 Evaluation Metrics
To quantify cross-lingual transferability, we follow the metric definitions introduced by Goldman et al. (2025). Let $\mathcal{L}$ denote the set of evaluation languages. For each ordered pair $(L_{\text{train}},L_{\text{test}})$ with $L_{\text{train}}\neq L_{\text{test}}$, model predictions can be categorized using the contingency matrix in Table 1.
@@ 4 Demonstrating LiveCLKTBench: A Case Study
We evaluate several open-source LLMs using a single knowledge injection strategy: Continual Pre-Training (Ke et al., 2023), where models continue pretraining directly on the target documents using the standard causal language modeling objective (next-token prediction) applied to unmodified text. Although this case study highlights Model Comparison, the framework is agnostic to the post-training method and also supports Strategy Comparison, enabling evaluation of different knowledge injection techniques for a single model.
@@
We consider instruction-tuned, open-source models in the 7–9B parameter range, including Gemma  [^4] https://huggingface.co/google/gemma-2-9b-it, Mistral  [^5] https://huggingface.co/mistralai/Ministral-8B-Instruct-2410666https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, Qwen  [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Llama  [^8] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, and OLMo  [^9] https://huggingface.co/allenai/OLMo-2-1124-7B-Instruct. These model families differ in pretraining data, tokenizer design, and alignment strategies, providing a diverse perspective on cross-lingual transfer.
@@ Appendix A Data Statistics
In LiveCLKTBench, the training set consists of raw text documents that contain information about each knowledge entity (e.g., a movie). Each training document is associated with multiple QA pairs that probe different factual aspects of the same entity (e.g., cast, storyline, release details). Since we evaluate cross-lingual transfer by post-training the model on the source-language documents, we split the corresponding QA pairs into validation (20%) and test subsets (80%). The validation subset is used to select the best checkpoint and to monitor learning progress during post-training. The test subset is held out and used only for final evaluation. Importantly, both validation and test questions are still grounded in the same knowledge document used for training and do not introduce any new information.
@@ Appendix B Data Quality Verification
To assess the reliability of our benchmark data, we recruited three annotators (two undergraduate students and one graduate student) to manually verify the quality of sampled question–answer pairs. Specifically, we randomly sampled 20% of entities from each domain and evaluated all QA pairs associated with them. Each question was independently annotated by three annotators based on the following criteria, and was considered a failure if it violated any of them.
  • F-1: The question is unrelated to the source document.
  • F-2: The question is relevant but overly general and not specific to the target entity.
  • F-3: The answer cannot be verified from the source document.
  • F-4: The answer is verifiable from the source document but is incorrect.
For annotation, we required full agreement among annotators. Questions with disagreement were discarded from the analysis. On the remaining items (98.8%), the validated QA accuracy (Pass Rate) is 95.2%. The invalid items include 3.6% cases of F-3 (from the movie domain) and 1.2% cases of F-4 (from the sports domain). These results suggest that our automated pipeline produces reasonably high-quality, fact-grounded QA pairs.